OpenStudy (anonymous):

Convert to the given unit. 24 ft to yd PLEASE HELP WILL GIVE MEDAL FOR BEST RESPONSE

OpenStudy (anonymous):

There are 3 feet in a yard, so 24 feet is 24/3 = 8 yards. There are 12 inches in a foot, so 70 inches is 60inches + 10inches = 12*5 inches + 10 inches = 5 feet, 10 inches.

OpenStudy (anonymous):

in other words it give you ?

OpenStudy (phi):

First you need to know how many feet in a yard (or how many yards to a foot) say you know 1 yard = 3 feet write this as a ratio or fraction with "3 feet" in the bottom. \[ \frac{ 1 \ yd}{3 \ ft} \] now multiply this fraction times 24 ft

OpenStudy (anonymous):

8

OpenStudy (anonymous):

like what @phi said

OpenStudy (anonymous):

@TH3UPR04R187 you have to give hints not answer thank you :)